Consider the following two stocks. Probabilities Stock ky) FA Recession1=35% 9% Normal 2=33% 8% Boom B=32% 15% Stock B 3%Using the correct answers from the previous questions, what is the expected return of the portfolio? Enter your answer as a p
Answer #1

As all the answer are computed and are correct. I'm calculating standard deviation of portfolio.

First, we will calculate the Co-variance of Stock A & B:-

Probability(P) Stock A (RA) P*RA Stock B (RB) P*RB​​​​​

Deviation of RA

(RA-ERA)

Deviation of RB

(RB-ERB)

Deviation of RA*Deviation of RB (Deviation of RA*Deviation of RB)*p
0.35 -9 -3.15 3 1.05 -13.29 -6.35 84.3915 29.537
0.33 8 2.64 -2 -0.66 3.71 -11.35 -42.1085 -13.8958
0.32 15 4.8 28 8.96 10.71 18.65 199.7415 63.917
ERA=4.29 ERB=9.35 COVAB= 79.5582

So, CovAB = 79.5582

-- Coefficient of Corrrelation(ρ) = CovABAσB

= 79.5582/(10.15*12.96)

= 0.60

So, now we will calculate Standard deviation of portfolio:-

\sigma_{P} =\sqrt{(W_{A})^{2}(\sigma _{A})^{2}+{(W_{B})^{2}(\sigma _{B})^{2}+2W_{A}W_{B}\sigma _{A}\sigma _{B}\rho _{AB}}

\sigma_{P} =\sqrt{(0.2)^{2}(10.15)^{2}+(0.8)^{2}(12.96)^{2}+2*0.2*0.8*10.15*12.96*0.6}

\sigma_{P} =\sqrt{4.1209+107.4954+25.2564}

\sigma_{P} =\sqrt{136.8727}

\sigma_{P} =11.70

So, Standard deviation of Portfolio is 11.70%

If you need any clarification regarding this solution, then you can ask in comments

If you like my answer then please Up-vote as it will be motivating.